If the compliment of an angle is one tenth the supplement of the angle than what is the measure of the angle? I am unsure where to start.....
Answer by pushpaharan(47)

let the angle be x.
therefore complement angle to x = 90-x
supplement angle = 180-x
but compliment of an angle is one tenth the supplement of the angle
90-x = 1/10(180-x)
10(90-x) = (180-x)
900-10x = 180- x
-10x + x = 180-900
-9x = -720
x=80
Hope you got it.:-)